the quantity of charge q (in coulombs) that has passed through a surface of area 2.00 cm^2 varies with time according to the equation q=4t^3 + 5t+6 , where t is in seconds.a)what is the instantaneous current through the surface at t=1.00 s? b)what is the value of the current density?
1
Expert's answer
2020-04-09T09:34:28-0400

a)


"I=\frac{dq}{dt}=4(3t^2)+5=12t^2+5"

"I(1)=12(1)^2+5=17\ A"

b)


"J(1)=\frac{I(1)}{A}=\frac{17}{2\cdot 10^{-4}}=85000\frac{A}{m^2}"